OpenStudy (anonymous):

What is the coordinate of point Q? A. 3.2 B. 3.5 C. 3.6 D. 3.7

OpenStudy (anonymous):

fan and medal but answers only!

OpenStudy (anonymous):

@YOUNGSUPERMAN

OpenStudy (anonymous):

@Ashes.boom

OpenStudy (anonymous):

@ashellyh

OpenStudy (anonymous):

@Rainbow_Dashie.

OpenStudy (anonymous):

@Allenismyname

OpenStudy (anonymous):

@huntin14

OpenStudy (anonymous):

There are 12 questions with enough help we can solve all!

OpenStudy (anonymous):

Please!

OpenStudy (ashes.boom):

is there a picture or a graph or something

OpenStudy (youngsuperman):

3.7 I think

OpenStudy (anonymous):

Thank you, now question 2

OpenStudy (anonymous):

Which is an example of an algebraic expression? A. d – 90 B. d < 7 C. 55 + v > 60 D. a + 5 = 72

OpenStudy (anonymous):

10 more once we answer this one

OpenStudy (anonymous):

6th grade math

OpenStudy (anonymous):

@iGreen

OpenStudy (ashes.boom):

i think its d

OpenStudy (anonymous):

thx

OpenStudy (youngsuperman):

I agree

OpenStudy (ashes.boom):

yea no problem

OpenStudy (anonymous):

Which statements are true according to the order of operations? Choose exactly two answers that are correct. A. 15 • (6 ÷ 3) = 90 ÷ 3 = 30 B. 4 + 12 ÷ 2 = 4 + 6 = 10 C. (20 – 5) • 3 = 15 • 3 = 45 D. 32 + 8 ÷ 4 = 40 ÷ 4 = 10

OpenStudy (anonymous):

9 more once this has been solved

OpenStudy (ashes.boom):

PEMDAS Parentheses Exponents Multiplication/division Addition/subtraction

OpenStudy (anonymous):

I know that from 5th grade I just need my two answers

OpenStudy (anonymous):

@EWack

OpenStudy (ashes.boom):

so if you know it then why don't you just solve the equation

OpenStudy (anonymous):

@Ashes.boom Sorry, It is just this is a little hard for me and I have to go somewhere soon so I wanted to get this done as soon as possible.

OpenStudy (ashes.boom):

okay okay well open study is being amazingly slow so i will inbox you okay

OpenStudy (anonymous):

okay, Thank you so much!

OpenStudy (anonymous):

Which statements are true according to the order of operations? Choose exactly two answers that are correct. A. 15 • (6 ÷ 3) = 90 ÷ 3 = 30 B. 4 + 12 ÷ 2 = 4 + 6 = 10 C. (20 – 5) • 3 = 15 • 3 = 45 D. 32 + 8 ÷ 4 = 40 ÷ 4 = 10

OpenStudy (anonymous):

Which options could be the next step in simplifying this expression? (18 + 23) + 7 Choose exactly two answers that are correct. A. 25 + 30 (use the distributive property) B. 30 + 25 (use the commutative property) C. 18 + 30 (use the associative property) D. 41 + 7 (use order of operations)

OpenStudy (anonymous):

@Rainbow_Dashie.

OpenStudy (anonymous):

are you still working on this? i was doing my classes

OpenStudy (anonymous):

Do you still need help? @TT04

OpenStudy (anonymous):

yes